OpenStudy (anonymous):

find the inverse of -7+ 5i

OpenStudy (anonymous):

The conjugate is -7-5i. If you mean inverse as: (-7+5i)^(-1) then you have: \[\frac{1}{-7+5i}*\frac{-7-5i}{-7-5i}=\frac{-7-5i}{74}\]